# ClutterGen: A Cluttered Scene Generator for Robot Learning

### v1

@@ 1 Introduction
Simulation has played an important role in advancing robot learning [1, 2, 3, 4] by providing a controlled yet versatile environment for developing and testing algorithms. Data-driven approaches, in particular, typically deploy robots into simulations to undergo extensive training across a variety of diverse and randomized settings to enable generalizable and adaptable behaviors. Significant advancements in robot learning have been achieved by randomizing object shapes [4, 5], textures [6, 7, 8, 9], and dynamics [10]. However, the layout of objects, despite being another critical factor, remains challenging to reach fully open-ended randomization.
Unlike object properties, which can be easily specified within a range without interfering with other objects, object layout must consider the presence of other objects and physical feasibility. For instance, arranging objects in a scene requires ensuring that they do not overlap and are placed in stable positions instead of falling down from the air. Existing efforts often prevent this issue by fixing the object bases [11, 4, 12, 13], but this strategy is not suitable for many objects like bottles or cups. As the number of objects increases within a limited space, generating a randomized yet stable object layout becomes exponentially difficult. Fig. 1(a) shows the challenge of using the widely adopted approach of random sampling and rejecting failure trails [14, 15, 16, 17] to generate valid scenes, with even seven objects on a table. Other methods require human manual specifications of object regions for local randomization [18, 19, 20] or apply discretization to the possible placement space to avoid collisions [15, 21, 22]. However, navigating and manipulating cluttered environments are essential challenges to deploying robot learning to the real world.
@@ 3.2 Cluttered Scene Generator
Observation and Action Space ClutterGen takes in the point cloud of the queried scene and object. If the current object placement fails, we inform the policy about its past actions and their impacts by providing a history of the queried object’s movement trajectories $M_{i}=\{m_{0},m_{1},...,m_{k}|m_{i}\in\mathbb{R}^{13}\},i=\{1,2,...,I\}$. Each $m$ includes the object’s position $\in\mathbb{R}^{3}$, orientation $\in\mathbb{R}^{4}$, linear velocity $\in\mathbb{R}^{3}$ and angular velocity $\in\mathbb{R}^{3}$ at each step. The hyperparameter $k$ determines the number of simulation steps to verify placement success, and $I$ is the predefined maximum number of attempts allowed.
@@ 4.1 Scene Generation Evaluation
Baselines Since most prior studies on scene generation rely on either pre-collected datasets or human-specified heuristics and constraints, we cannot directly compare with them. The closet method, which is also widely adopted [14, 15, 16, 17] in recent literature, is the random rejection sampling (RSS) algorithm. We also compare our methods with other alternative settings.
@@ Acknowledgments
This work is supported by ARL STRONG program under awards W911NF2320182 and W911NF2220113, and by DARPA FoundSci program under award HR00112490372.
@@
Selection of Maximum Allowed Attempts. We test the performance of ClutterGen with different maximum attempts on group 1 objects. Increasing allowed attempts will boost the performance. However, too many attempts will slow down the training process and focus on only place one object. We set 5 attempts at most to balance the performance and training speed.

### v2

@@ 1 Introduction
Simulation has played an important role in advancing robot learning [1, 2, 3, 4, 5] by providing a controlled yet versatile environment for developing and testing algorithms. Data-driven approaches, in particular, typically deploy robots into simulations to undergo extensive training across a variety of diverse and randomized settings to enable generalizable and adaptable behaviors. Significant advancements in robot learning have been achieved by randomizing object shapes [4, 6, 7], textures [8, 9, 10, 11], and dynamics [12]. However, the layout of objects, despite being another critical factor, remains challenging to reach fully open-ended randomization.
Unlike object properties, which can be easily specified within a range without interfering with other objects, object layout must consider the presence of other objects and physical feasibility. For instance, arranging objects in a scene requires ensuring that they do not overlap and are placed in stable positions instead of falling down from the air. Existing efforts often prevent this issue by fixing the object bases [13, 4, 14, 15], but this strategy is not suitable for many objects like bottles or cups. As the number of objects increases within a limited space, generating a randomized yet stable object layout becomes exponentially difficult. Fig. 1(a) shows the challenge of using the widely adopted approach of random sampling and rejecting failure trials [16, 17, 18, 19] to generate valid scenes, with even seven objects on a table. Other methods require human manual specifications of object regions for local randomization [20, 21, 22] or apply discretization to the possible placement space to avoid collisions [17, 23, 24]. However, navigating and manipulating cluttered environments are essential challenges to deploying robot learning to the real world.
@@ 3.2 Cluttered Scene Generator
Observation and Action Space ClutterGen takes in the point cloud of the queried scene and object. If the current object placement fails, we inform the policy about its past actions and their impacts by providing a history of the queried object’s movement trajectories $M_{i}=\{m_{0},m_{1},...,m_{k}|m_{i}\in\mathbb{R}^{13}\},i=\{1,2,...,I\}$. Each $m$ includes the object’s position $\in\mathbb{R}^{3}$, orientation quaternion $\mathbb{H}\in\mathbb{R}^{4}$, linear velocity $\in\mathbb{R}^{3}$ and angular velocity $\in\mathbb{R}^{3}$ at each step. The hyperparameter $k$ determines the number of simulation steps to verify placement success, and $I$ is the predefined maximum number of attempts allowed.
@@ 4.1 Scene Generation Evaluation
Baselines Since most prior studies on scene generation rely on either pre-collected datasets or human-specified heuristics and constraints, we cannot directly compare with them. The closest method, which is also widely adopted [16, 17, 18, 19] in recent literature, is the random rejection sampling (RRS) algorithm. We also compare our methods with other alternative settings.
@@ Acknowledgments
This work is supported by ARL STRONG program under awards W911NF2320182 and W911NF2220113, by DARPA FoundSci program under award HR00112490372, and DARPA TIAMAT program under award HR00112490419.
@@
Selection of Maximum Allowed Attempts. We test the performance of ClutterGen with different maximum attempts on group 1 objects. Increasing allowed attempts will boost the performance. However, too many attempts will slow down the training process. We set 5 attempts at most to balance the performance and training speed.